INAF LBC+MODS queue observing, 2011 October 20-21

Observer: K. Boutsia, M. Fumana, P. Santini
Telescope Operator: G. Bechetti
Support Astronomer: D. Thompson
Instrument: LBC + MODS

Summary

Clear skies at the beginning of the night. Most probable photometric.
Variable seeing at the beginning of the night (up to 1.7) but for most of the night the seeing was <1.0''.
During flats there has been a mirror ventilation alarm that did not influence observations.
While switching from LBC to MODS we had a fault at the LDG rotator and 1hr30min have been lost.

Data Summary

Project Objects Filter ExpTime Notes
Bragaglia NGC2141 Bbess - Vbess / Ibess 20min the seeing was <0.8''
Clementini AndXXV Bbess - Vbess 2hr 20min 1 or 2 OBs at the beginning of the night are out of spec with seeing 1.5
Clementini AndXXVII Bbess - Vbess 2hr 20min 1 or 2 OBs at the beginning of the night are out of spec with seeing 1.5
STD Stetson_L92 Bbess - Vbess / Ibess   1 position - airmass 1.2 and 1.46
STD SA113 Bbess - Vbess / Ibess   4 positions - airmass 1.25
STD SA98 Bbess - Vbess / Ibess   4 positions - airmass 1.9
Donnarruma MAC0717b2 r_sdss 1hr15min the seeing was <1.0 - STD, flats and arcs for this mask can be found in UT 20112022

Details

12:55 Starting mirror ventilation is causing alarm... They go up to check.
  • we have a bad motor

01:03 Test SkyFlats
  • BV 0 (1.5sec-1sec) - scale 1.2 - B 16000count / V 12000 counts (01:10)
  • BV_180 (1.5sec - 1sec) scale 2.5 counts too low - stopped
  • BI_180 (3sec-1sec) scale 5 - B 20000count (a lot of stars) - I 30000counts (stars)

01:19 Slew to pointig star WT10_219
  • dofpia,/first
  • stopped after the 4th iteration - it is almost converging

01:34 lbcrangebal on WT10_219

01:40 Slew to SA113 for standard calibration (elev~53 , airmass=1.25)
  • dither in 4 positions

01:47 Slew to M31_AndXXV_400s_5
  • dofpia

01:56 Starting M31_AndXXV_400s_5 obs
  • seeing seems to be higher than 20/10/2011
  • dimm seeing 1.8
  • seeing on first images ~1.35
  • seeing on second images: ~1.60 (blue), ~1.88 (red)

02:19 Slew to M31_AndXXVII_400s_5 obs
  • dofpia

02:23 Starting M31_AndXXVII_400s_5 obs
  • dimm seeing 2.7
  • seeing on first images ~1.75 (blue), ~1.80 (red)
  • seeing on images ~1.65 (blue), ~1.77 (red)

02:46 Slew to M31_AndXXV_400s_6
  • dofpia
  • temperature is dropping 1 degree per hour, this is probably causing the variable seeing

02:55 M31_AndXXV_400s_6.ob
  • dimm seeing 1.6
  • seeing on images ~1.2 (blue), <1.1 (red)
  • seeing on last images ~1.2 (blue), <1.2 (red)

03:19 Slew to M31_AndXXVII_400s_6
  • dofpia

03:22 M31_AndXXVII_400s_6.ob
  • seeing on first images ~1. 0 (blue), <1. 0 (red)
  • seeing on images <1.0 (blue), <1.0 (red)
  • seeing on last images ~1. 1 (blue), ~1. (red)

03:46 Slew to M31_AndXXV_400s_7

03:50 M31_AndXXV_400s_7.ob
  • seeing measured on images varies from 1.1'' to 1.35''

04:13 Slew to M31_AndXXVII_400s_7
  • dofpia

04:17 M31_AndXXVII_400s_7.ob
  • seeing measured on images varies from ~0.88(blue) ~0.84 (red)

04:39 Slew to M31_AndXXV_400s_8
  • dimm seeing ~1.1-1.3
  • dofpia

04:43 M31_AndXXV_400s_8.ob
  • seeing measured on images varies from ~0.86 (blue) ~1 (red)

05:07 Slew to M31_AndXXVII_400s_8
  • dofpia

05:14: M31_AndXXVII_400s_8.ob
  • dimm seeing <1
  • seeing measured on first images ~0.8 (blue) ~0.9 (red)
  • seeing measured on second images ~0.86 (blue) <1 (red)
  • seeing measured on last images ~0.86 (blue) <0.95 (red)

05:37 Slew to Stetson_L92_1 for standard calibration (elev~56.8 , airmass=1.2)

05:44 Slew to M31_AndXXV_400s_9
  • dofpia

05:52 M31_AndXXV_400s_9.ob
  • seeing measured on first images ~0.7 (blue) <1 (red)
  • seeing measured on second images ~0.74 (blue) ~0.77 (red)
  • seeing measured on last images ~0.72 (blue) <0.86 (red)

06:16 Slew to M31_AndXXVII_400s_9
  • dofpia

06:24 M31_AndXXVII_400s_9.ob
  • seeing measured on first images <0.74 (blue) 0.74-0.97 (red)
  • seeing measured on second images ~0.72 (blue) ~0.72-0.83 (red)
  • seeing measured on last images <0.7 (blue) <0.83 (red)

06:47 Slew to M31_AndXXV_400s_10
  • dofpia

06:51 M31_AndXXV_400s_10.ob
  • seeing measured on first images 0.79 (blue) <0.83 (red)
  • seeing measured on second images ~0.8 (blue) <0.83 (red)
  • seeing measured on last images 0.85 (blue) ~0.83 (red)

07:14 Slew to M31_AndXXVII_400s_10
  • dofpia

07:21 M31_AndXXVII_400s_10.ob
  • seeing measured on first images ~0.72 (blue) <0.79 (red)
  • seeing measured on second images ~0.76

07:43 Slew to M31_AndXXV_400s_11
  • dofpia

07:48 M31_AndXXV_400s_11.ob
  • seeing measured on first images ~0.77 (blue) 0.72-0.81 (red)
  • seeing measured on second images ~0.72 (blue) ~0.7 (red)
  • seeing measured on last images 0.67

08:12 M31_AndXXVII_400s_11.ob
  • seeing measured on first images ~0.79 (blue) <0.79 (red)
  • seeing measured on second images <0.76 (blue) ~0.67 (red)
  • seeing measured on last images ~0.79 (blue) ~0.67 (red)

08:35 Slew to Stetson_L92_3 for standard calibration (airmass=1.46)

08:44 Slew to SA98_B for standard calibration (airmass=1.9)

08:52 Slew to ngc2141
  • dofpia

08:56 ngc2141.ob
  • seeing measured on images <1

09:11 SWITCH TO MODS

09.24 one rotator is stuck and MODS can not be positioned

10:06 motor problem solved - 40min time lost for telescope.

10:11 preseting to ACT0346 for pointing correction with MODS
  • dimm is around 1.0

10:24 preset to MAC0717 (mods.7189.201130b2.mms) - Donnarumma
  • two presets failed - we don't know why.
  • the rotator faulted again - this is why we are off source and the preset failed.

10:45 the rotator problem seems solved

10:59 send preset again

11:04 the throughslit image has arrived and we run modsAlign
  • offsetpointing -0.222 1.78 -0.254 detxy rel

11:20 execMODS mods.7189.201130b2. obs
  • mods1r.20111021.0004.fits 2250sec
  • mods1r.20111021.0004.fits 2250sec

12:37 - acqMODS standard_star.acq
  • preset failed - repeat
  • Dave adjusted astigmatism by hand so that we don't loose the guide star again
  • the star is well centered in the slit after the acquisition, no need to align the slit with modsAlign

12:43 execMODS standard_star.obs
  • the first image of the standard is saturated.
  • we try again with 30 sec exposure - we get a max of 20000 - it should be fine - seeing today is better than yesterday
  • mods1r.20111021.0007 - 0008 - 0009.fits

12:54 Closing telescope - moving in MODS calibration UNIT
  • flats and arcs for MOS mask (actually the mask for the arcs is the b1 - for b2 arcs look at 20111022)
  • flats and arcs for slit (used for the standard)
  • slitless flats
  • allbias script
Topic revision: r12 - 26 Nov 2011, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback